OpenStudy (anonymous):

Suzie worked the following problem incorrectly. Part 1) In which line was the mistake made? Part 2) Show your work and give the correct answer? Line 1 50 ÷ 5(3 − 1) − 4 Line 2 50 ÷ 5(2) − 4 Line 3 50 ÷ 10 − 4 Line 4 5 – 4 Line 5 1

OpenStudy (anonymous):

\[50\div5\times 2-4=10\times 2-4\]

OpenStudy (anonymous):

The whole thing is right? I am lost on this question /:

OpenStudy (tiffanymak1996):

wait, satellite is right!

OpenStudy (tiffanymak1996):

well, and that's the only solution to the question anyway...

OpenStudy (anonymous):

): I feel like I am in china with this. It's so confusing. Can you guys explain more by possible chance?

OpenStudy (tiffanymak1996):

sry, that i mislead you...

OpenStudy (anonymous):

\[50\div5\times (3-1)-4\] parentheses first \[50\div 5\times 2-4\] then multiplication and division AS YOU READ FROM LEFT TO RIGHT \[10\times 2-4\] \[20-4\] \[16\]

OpenStudy (anonymous):

the point is, you do not do the multiplication first, then the division you do them as you read from left to right, so if the division is first, you do that first

OpenStudy (tiffanymak1996):

how the question portrays the lines, it's specifically designed to trick you into thinking that what it supplied make sense, what you should do is to separate everything so that it'll be clearer.

OpenStudy (anonymous):

here is an example \[100\div 10\times 2=10\times 2=20\] NOT \(100 \div 20=5\)

OpenStudy (anonymous):

Ok.. I think I might understand now.. which line is the mistake?

OpenStudy (tiffanymak1996):

line 3

OpenStudy (tiffanymak1996):

50 ÷ 10 − 4 the 10 is incorrect

OpenStudy (anonymous):

Part 1. The mistake happened in line 3. Part 2. 50÷5×(3−1)−4 parentheses first 50÷5×2−4 then multiplication and division AS YOU READ FROM LEFT TO RIGHT 10×2−4 20−4 16

OpenStudy (anonymous):

That is what I have for the answer to the question..

OpenStudy (tiffanymak1996):

yeah, your right.

OpenStudy (anonymous):

are you sure you said the correct answer was 10 ...

OpenStudy (tiffanymak1996):

no, what i meant was that 10 was the mistake in line 3

OpenStudy (anonymous):

Oh okay. I understand now :) Thank you so much. hhaha.

OpenStudy (tiffanymak1996):

:)

OpenStudy (anonymous):

A concession stand charges $1.50 per soda and $2.00 per pretzel. Write an expression for the amount of money you would need to buy x sodas and y pretzels.

OpenStudy (anonymous):

I also have that one if you can help me with it if not it's fine : )